OpenStudy (anonymous):

Part A.)Simplify 4(3x+2)-2, explain each step and identify the property used. (show all steps) (Part B.) Use the distributive property used to rewrite your expression for part A as a product

OpenStudy (anonymous):

Part A: Step 1: Distributive property 4(3x+2)−2 4×3x=12x 4×2=8 ------------------------------------------ 12x+8−2 Subtract 8−2=6 ------------------------------------------ 12x+6 < ---- answer Part B: I don't understand